OpenStudy (anonymous):

Help please stuck. 1. Write an explicit formula for the sequence; 3,-1,-5,-9,-13,... then find a14

hartnn (hartnn):

do u see any pattern in that sequence ?

OpenStudy (anonymous):

umm im not sure of what it is :/

OpenStudy (fibonaccichick666):

try and look at the differnces between each term

OpenStudy (anonymous):

so -4

OpenStudy (fibonaccichick666):

ok so now, if there is a difference of -4 between each term, how could you represent that as an eq?

OpenStudy (anonymous):

3+(n-1)-4?

OpenStudy (fibonaccichick666):

without the correct notation that is gibberish^ please use the full notation

OpenStudy (anonymous):

an=a1+(n-1) d an=3+(n-1)-4?

OpenStudy (fibonaccichick666):

assume n starts at 1, check it. Does it give you 3? at n=2 do you get -1?

OpenStudy (anonymous):

sorry if this sounds dumb but would this be the correct way to set it up?

OpenStudy (fibonaccichick666):

good try, but not quite

OpenStudy (fibonaccichick666):

so your n is going to start at 1

OpenStudy (fibonaccichick666):

let me find you an instructional video

OpenStudy (fibonaccichick666):

^Watch that^

OpenStudy (anonymous):

 Thanks that helped.. would this be correct?

OpenStudy (fibonaccichick666):

uhm not quite, so here is an example for the summation \(b_n\) up to term 50

OpenStudy (fibonaccichick666):

\[\huge\Sigma^{51}_{n=1} b_{n-1}\]

OpenStudy (fibonaccichick666):

help at all?

OpenStudy (fibonaccichick666):

also maybe read through here

OpenStudy (fibonaccichick666):

http://openstudy.com/study#/updates/5361d2f5e4b033b553834456 it's how someone else figured out similar

OpenStudy (anonymous):

Okay so am i close? \[\sum_{n=1}^{14}3-4n\]

OpenStudy (fibonaccichick666):

very close

OpenStudy (fibonaccichick666):

so your first term should be 3, if n=1, do you get 3?

OpenStudy (anonymous):

yes. is the number ontop incorrect (14)?

OpenStudy (fibonaccichick666):

check again, 3-4(1)=?

OpenStudy (anonymous):

wait i got it so it would be 1-4(1)

OpenStudy (fibonaccichick666):

o.O? Not sure how you got that

OpenStudy (fibonaccichick666):

i was just asking for what that equaled

OpenStudy (anonymous):

youre right so n=0

OpenStudy (fibonaccichick666):

3-4(1)=? <----what is the question mark?

OpenStudy (fibonaccichick666):

yes, n can indeed =0

OpenStudy (anonymous):

I'm sorry i changed the numbers and got mixed up.

OpenStudy (anonymous):

because 3-4(1)=-1

OpenStudy (fibonaccichick666):

if you change that index, what does the top number have to become?

OpenStudy (anonymous):

would it just be -12?

OpenStudy (fibonaccichick666):

it's never negative, nor is it 12 either, how did you get that?

OpenStudy (anonymous):

Because I thought I had to minus -1 from the -13.

OpenStudy (fibonaccichick666):

well, one -13? why?

OpenStudy (anonymous):

Because in the other example I saw that it went down one number

OpenStudy (fibonaccichick666):

, but why negative 13?

OpenStudy (anonymous):

Oh because of the last term in the sequence

OpenStudy (fibonaccichick666):

ok, yea, the number on top is never negative

OpenStudy (anonymous):

Okay I see. So would it be 13? or

OpenStudy (fibonaccichick666):

yes

OpenStudy (anonymous):

\[\sum_{n=0}^{13}3-4n\]

OpenStudy (anonymous):

thanks so much. youre the best

OpenStudy (anonymous):

finally understood this

OpenStudy (fibonaccichick666):

np happy you get it, remember, you can ALWAYS check your answer just by finding the first few terms

OpenStudy (fibonaccichick666):

see if they match up, if they do, you're good

OpenStudy (anonymous):

I have a question, when it says find a14 what do i have to do?

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

can you tell me what term a_1 is?

OpenStudy (anonymous):

its the same equation as 3,-1,-5,-9,-13,... then find a14

OpenStudy (anonymous):

so a1 is 3

OpenStudy (fibonaccichick666):

mhmm

OpenStudy (fibonaccichick666):

and how do you get that in your eq?

OpenStudy (fibonaccichick666):

it occurs at n=?

OpenStudy (anonymous):

0?

OpenStudy (fibonaccichick666):

mhmm

OpenStudy (anonymous):

so then how would i find a14?

OpenStudy (fibonaccichick666):

if to find a1 you put in n=0, to find a14 you put in n=?

OpenStudy (anonymous):

I dont know if im wrong but would it be n=7

OpenStudy (fibonaccichick666):

how did you get that? no, just follow the logic from above what is the difference between 0 and 1?

OpenStudy (anonymous):

1